# Entropy conservative discretization of compressible Euler equations with an arbitrary equation of state

### v1

@@ 5 Entropy fluxes
It is known that the scheme of Ranocha, specified by the fluxes in Eq. (21), is in locally-conservative form for the entropy equation, i.e. it admits induced entropy fluxes. In fact, at p. 26 of [42] (which is an extended version of [41]), the entropy numerical flux consistent with the formulation expressed in Eq. (21) is reported, and can be written, in our notation, as:
@@
The scheme of Ranocha has been derived in Sec. 4.2 as a special case, for ideal gases, of the novel and more general formulation presented in Sec. 4.1 for gases with an arbitrary equation of state. However, this last formulation has been obtained by enforcing global conservation of entropy, which does not imply, in general, that a local entropy flux exists. It is then legitimate to ask if the local conservation of entropy present in the ideal gas formulation is also present in the more general case. To this aim, it is useful to recall here some results contained in a recent paper by Coppola and Veldman [16], where, in the context of an abstract theoretical analysis on the relations between the concepts of global and local conservation of primary and secondary invariants, a quite general equivalence between global and local conservation is estabilished, at least in the framework of classical FD discretizations. This result is used here to show that our formulation, obtained by imposing global conservation of entropy, admits local entropy fluxes for arbitrary gas models. Of course, these fluxes are consistent with the ones already known for the special case of ideal gases.
@@
$\begin{array}[]{c c c}\int\mathcal{C}_{\rho\phi}\,\text{d}x=0&\quad\longrightarrow&\mathbf{1}^{\text{T}}{\sf C}{\sf v}=0,\end{array}$ (29)
@@
$\begin{array}[]{c c c}\mathbf{1}^{\text{T}}{\sf C}=\mathbf{0}^{\text{T}}&\quad\iff&{\sf C}{\sf v}={\sf f}-\widetilde{{\sf f}}.\end{array}$ (30)
Moreover, the theorem reported in the Appendix A1 of [16] gives also an explicit formula for the derivation of the numerical flux ${\sf f}$. We will not go into the details of this derivation and refer the interested reader to the original paper and in particular to its Appendix A. Here we use this result to infer the existence of numerical entropy fluxes in all cases and obtain them with a more direct argument based on the derivation made in Sec. 4.1.
@@ 6 High-order formulation
In deriving the EC internal-energy flux in Eq. (20), we implicitly assumed that the interpolations used for the mass flux $\mathcal{F}_{\rho}$ and for the velocity in the pressure term $\mathcal{P}_{\rho e}$ are two-point averages of the variables at the nodes $i$ and $i+1$ adjacent to the face $i+1/2$, thus implying a second-order accuracy for the whole procedure. However, the treatment can be extended to higher-order formulations by using a result contained in the Appendix A of the recent paper by De Michele and Coppola [29], which we recall here for convenience.
@@
$02\sum_{k=1}^{L}a_{k}\left[\sum_{i}\left(\mathcal{F}^{i,k}_{\rho e}\delta^{\,+}_{k}\dfrac{1}{T}-\mathcal{F}^{i,k}_{\rho}\delta^{\,+}_{k}\dfrac{g}{T}+\overline{u}^{k}\delta^{\,+}_{k}\dfrac{p}{T}\right)\right]=$ 0 (49)
for which the validity of Eq. (44) guarantees that each contribution in the inner parentheses vanishes. This shows that if the high-order internal-energy flux is built according to Eq. (40) with $\mathcal{F}^{i,k}$ specified as in Eq. (44), the global production of entropy is nullified, as in the second order case.
@@ 7 Numerical results
In this section, the results of numerical simulations of the compressible Euler equations are presented to verify the theoretical analysis and to assess the effectiveness of the newly proposed procedure. They are carried out for two cubic equations of state (Van der Waals and Peng-Robinson models) and for two configurations, a one-dimensional traveling density-wave test and a two-dimensional transcritical double shear layer.
The balance equations for mass, momentum and total energy (1)–(3) have been discretized by using a standard KEP formulation for mass and momentum, whereas the newly derived EC flux given in Eq. (20) has been used for the internal-energy contribution in the total-energy flux, resulting in the set of fluxes:
@@
where $\mathcal{F}_{\rho e}^{EC}$ is the flux in Eq. (20). The formulation in Eq. (50) is second-order accurate; its extension to higher-order cases is made according to the theory in Sec. 6. All the simulations have been carried out with a fourth-order explicit Runge-Kutta scheme for time integration, with Courant number taken sufficiently small in order to nullify the temporal errors and focus only on the spatial discretization errors.
As regards the singularity in the flux $\mathcal{F}_{\rho e}^{EC}$ defined in Eq. (20), in all the tests we implement a simple fix consisting in the local use of a non-singular flux in regions in which the temperature distribution is nearly uniform. In particular, the implementation is made by monitoring the local values of $\delta^{\,+}T$. If this quantity is below a certain tolerance, we locally switch to a non-singular numerical flux. For this purpose, we use the AEC schemes presented by De Michele and Coppola [29], which are designed to have excellent EC properties for ideal gases. This class of schemes, which are nonsingular and based on economic algebraic operations, constitutes a hierarchy of numerical fluxes based on a suitable truncated series expansion. They are designed to satisfy the KEP and PEP properties at each finite truncation, and asymptotically preserve the entropy, in the sense that the numerical error on global entropy production (in the ideal gas case) can be arbitrarily reduced by including additional terms in the expansion. In fact, taking two terms in the series expansion has been verified to be enough to avoid singularities without spoiling the EC properties of the new scheme.
@@
[…]
where the ideal-gas behaviour, that is represented by the integral contributions, is given by the 5-coefficient polynomial fitting for $c_{p}(T)$ reported in [46].
@@
which can be seen as the simplest modification of the ideal-gas equation of state, taking into account the effective volume of molecules ($b$) and the interaction between the molecules themselves via the potential-energy related term $a/\tilde{v}^{2}$. Despite its simplicity, such model remains widely used for highly-transcritical fluid flow simulations [47]. On the other hand, the Peng-Robinson model [48] is
@@
Figure 1: Convergence analysis for high-order formulations. (a) Global error on internal energy as a function of the mesh size. Red lines with filled circles refer to the Van der Waals equation of state, black lines with empty circles refer to the Peng-Robinson model. (b) Entropy conservation error for $2$nd, $4$th, $6$th and $8$th order (from top to bottom) on a uniform mesh of $N=34$ intervals.
@@
The density-wave problem is a classical test case for compressible Euler equations, typically used to assess the PEP property. In our framework, however, the density-wave test has been chosen as a simple one-dimensional case in which the entropy-conservation properties of the novel formulation can be assessed with minimum occurrences of locally uniform temperature distributions. This allows us to test the original fluxes by minimizing the use of the alternative non-singular formulations.
@@
[…]
@@
Fig. 2 shows the results of simulations with final time $t=0.6$ s and $N=40$ intervals, with a Courant number set to $0.005$ for the Van der Waals (2nd order, Fig. 2(a)) case and $0.002$ for the Peng-Robinson (4th order, Fig. 2(b)) case, in order to be able to neglect time-integration errors. The performances of the novel fluxes are compared with those of the Ranocha’s and AEC(0) formulations. The insets clearly show that the new formulation reduces the global error on entropy production up to the machine precision, whereas the classical formulations exhibit a non-negligible spurious production of entropy which accumulates in time. As it is possible to note from Fig. 2, the AEC(0) scheme seems to perform better than the Ranocha scheme in this particular case. The same simulations, not included here, have also been carried out with the AEC scheme with two additional terms of the series expansion which has been chosen as the non-singular flux to be used locally in the presence of singularities for the new EC flux. The results showed an essentially identical behaviour to that of the scheme of Ranocha regarding the error $\langle\rho s\rangle$, which is in agreement with the theory presented in [29].
@@ 7.2 Transcritical double shear layer
A two-dimensional double shear layer is simulated in the rectangle $[-L_{x},L_{x}]\times[-L_{y},L_{y}]$, with $L_{x}=0.5\,\text{m}$ and $L_{y}=0.25\,\text{m}$, discretized in $N_{x}\times N_{y}=32\times 16$ intervals on a uniform mesh with periodic boundary conditions and initial conditions
@@
[…]
@@
From Fig. 3, reporting the entropy production for the various schemes, it is seen that the theoretical conservation properties of the new formulation are confirmed also in this test. Again, the harmonic numerical flux on internal energy provides a smaller error on the conservation of $\rho s$ if compared to the Ranocha scheme, although, by looking at the results in Fig. 2 too, the difference between the two schemes seem to be reduced by the higher-order formulation.
@@ 8 Conclusions
The derived formulation is, to the best of our knowledge, the first entropy-conserving discretization of the compressible Euler equations specifically designed for real gases. Its structure-preserving properties make it an ideal candidate for high-fidelity simulations of turbulent flows where an arbitrary equation of state is used. Moreover, it constitutes the first building block toward a fully structure-preserving discretization of compressible flow equations for real gas applications. In fact, the condition introduced in this work on the discretization of the convective term of internal energy introduces a whole family of EC schemes for which, however, no constraint is imposed on the choice of the numerical mass flux. The use of this addition degree of freedom could be exploited to impart the scheme with supplementary structure-preserving properties, an example of which being the pressure equilibrium preservation property that is often desirable in the contexts of multi-component gases or multi-phase flows. The exploration of this topic is left for future research.

### v2

@@ 5 Entropy fluxes
It is known that the scheme of Ranocha, specified by the fluxes in Eq. (21), is in locally-conservative form for the entropy equation, i.e. it admits induced entropy fluxes. In fact, at p. 26 of [48] (which is an extended version of [25]), the entropy numerical flux consistent with the formulation expressed in Eq. (21) is reported, and can be written, in our notation, as
@@
The scheme of Ranocha has been derived in Sec. 4.2 as a special case, for ideal gases, of the novel and more general formulation presented in Sec. 4.1 for gases with an arbitrary equation of state. However, this last formulation has been obtained by enforcing global conservation of entropy, which does not imply, in general, that a local entropy flux exists. It is then legitimate to ask if the local conservation of entropy present in the ideal gas formulation is also present in the more general case. To this aim, it is useful to recall here some results contained in a recent paper by Coppola and Veldman [16], where, in the context of an abstract theoretical analysis on the relations between the concepts of global and local conservation of primary and secondary invariants, a quite general equivalence between global and local conservation is established, at least in the framework of classical FD discretizations. This result is used here to show that our formulation, obtained by imposing global conservation of entropy, admits local entropy fluxes for arbitrary gas models. Of course, these fluxes are consistent with the ones already known for the special case of ideal gases.
@@
$\begin{array}[]{c c c}\int\mathcal{C}_{\rho\phi}\,\text{d}x=0&\quad\longrightarrow\quad&\mathbf{1}^{\text{T}}{\sf C}{\sf v}=0,\end{array}$ (29)
@@
[…]
Moreover, the theorem reported in [16, Appendix A1] gives also an explicit formula for the derivation of the numerical flux $\sf f$}. We will not go into the details of this derivation and refer the interested reader to the original paper and in particular to its Appendix A. Here we use this result to infer the existence of numerical entropy fluxes in all cases and obtain them with a more direct argument based on the derivation made in Sec. 4.1.
@@ 6 High-order formulation
In deriving the EC internal-energy flux in Eq. (20), we implicitly assumed that the interpolations used for the mass flux $\mathcal{F}_{\rho}$ and for the velocity in the pressure term $\mathcal{P}_{\rho e}$ are two-point averages of the variables at the nodes $i$ and $i+1$ adjacent to the face $i+1/2$, thus implying a second-order accuracy for the whole procedure. However, the treatment can be extended to higher-order formulations by combining two-point fluxes of variable width with suitable coefficients. This approach has been used several times in the context of compressible Euler equations. LeFloch et al. [50] considered the two-point EC flux studied by Tadmor in a FV formulation, whereas Pirozzoli [10] used a similar framework to extend to high order the bilinear fluxes used in FD contexts. Fisher and Carpenter [51] treated the case of non-periodic boundary conditions for the Tadmor flux, and Ranocha [25] extended it to more general flux functions. We follow an adaptation of this last treatment contained in the Appendix A of the recent paper by De Michele and Coppola [30], which we recall here for convenience.
@@
$2\sum_{k=1}^{L}a_{k}\left[\sum_{i}\left(\mathcal{F}^{i,k}_{\rho e}\delta^{\,+}_{k}\dfrac{1}{T}-\mathcal{F}^{i,k}_{\rho}\delta^{\,+}_{k}\dfrac{g}{T}+\overline{u}^{k}\delta^{\,+}_{k}\dfrac{p}{T}\right)\right]=0$ (49)
for which the validity of Eq. (44) guarantees that each contribution in the inner parentheses vanishes. This shows that if the high-order internal-energy flux is built according to Eq. (40) with $\mathcal{F}^{i,k}$ specified as in Eq. (44), the global production of entropy is nullified, as in the second-order case.
@@ 7 Numerical results
In this section, the results of numerical simulations of the compressible Euler equations are presented to verify the theoretical analysis and to assess the effectiveness of the newly proposed procedure. They are carried out for two cubic equations of state (Van der Waals and Peng-Robinson models) and for three configurations: a one-dimensional traveling density-wave test, a two-dimensional transcritical double shear layer, and the three-dimensional Taylor-Green vortex in supercritical conditions. The balance equations for mass, momentum and total energy (1)–(3) have been discretized by using a standard KEP formulation for mass and momentum, whereas the newly derived EC flux given in Eq. (20) has been used for the internal-energy contribution in the total-energy flux, resulting in the set of fluxes:
@@
where $\mathcal{F}_{\rho e}^{EC}$ is the flux in Eq. (20). The formulation in Eq. (50) is second-order accurate; its extension to higher-order cases is made according to the theory in Sec. 6. All the simulations have been carried out with a fourth-order explicit Runge-Kutta scheme for time integration, except for the three-dimensional tests in Sec. 7.3, where a third-order, low-storage Runge-Kutta method is used. The Courant number is taken sufficiently small in order to nullify the temporal errors on entropy conservation and focus solely on the spatial discretization errors. Of course, higher values of the Courant number can be used when such a specific need is not required.
As regards the singularity in the flux $\mathcal{F}_{\rho e}^{EC}$ defined in Eq. (20), in all the tests we implement a simple fix consisting in the local use of a non-singular flux in regions in which the temperature distribution is nearly uniform. In particular, the implementation is made by monitoring the local values of $\delta^{\,+}T$. If this quantity is below a certain tolerance, we locally switch to a non-singular numerical flux. The threshold for the local values of $\delta^{\,+}T$ is, at present, empirical and test-based, and will be declared for each numerical test shown. We found that a relatively high threshold value for $\delta^{\,+}T$ does not spoil the EC property enforced by the novel formulation, whereas too small values could lead to losing precision in the evaluation of the fractions in (20), with the appearance of jumps in the temporal evolution of the global entropy-conservation error. The non-singular numerical fluxes used as a fix in this paper are the AEC schemes presented by De Michele and Coppola [30], which are designed to have excellent EC properties for ideal gases. This class of schemes, which are nonsingular and based on economic algebraic operations, constitutes a hierarchy of numerical fluxes based on a suitable truncated series expansion. They are designed to satisfy the KEP and PEP properties at each finite truncation, and asymptotically preserve the entropy, in the sense that the numerical error on global entropy production (in the ideal gas case) can be arbitrarily reduced by including additional terms in the expansion. In fact, taking two terms in the series expansion has been verified to be enough to avoid singularities without spoiling the EC properties of the new scheme.
@@
$\Delta e^{RG}_{1\rightarrow 2}$ $=-(e^{RG}-e^{IG})_{T_{1},p_{1}}+\int_{T_{1}}^{T_{2}}c_{v}(T)\,dT+(e^{RG}-e^{IG})_{T_{2},p_{2}}$ (51)
$\Delta s^{RG}_{1\rightarrow 2}$ $=-(s^{RG}-s^{IG})_{T_{1},p_{1}}+\int_{T_{1}}^{T_{2}}\frac{c_{p}(T)}{T}\,dT-R_{0}\log{\frac{p_{2}}{p_{1}}}+(s^{RG}-s^{IG})_{T_{2},p_{2}}$ (52)
where the ideal-gas behavior, that is represented by the integral contributions, is given by the 5-coefficient polynomial fitting for $c_{p}(T)$ reported in [53].
@@
which can be seen as the simplest modification of the ideal-gas equation of state, taking into account the effective volume of molecules ($b$) and the interaction between the molecules themselves via the potential-energy-related term $a/\tilde{v}^{2}$. Despite its simplicity, such model remains widely used for highly-transcritical fluid flow simulations [54]. On the other hand, the Peng-Robinson model [55] is
@@
Figure 1: Convergence analysis for high-order formulations. (a) Global error on internal energy as a function of the mesh size. Red lines with filled circles refer to the Van der Waals equation of state, black lines with empty circles refer to the Peng-Robinson model. (b) Entropy conservation error for 22nd, 44th, 66th and 88th order (from top to bottom) on a uniform mesh of $N=34$ intervals.
@@
The density-wave problem is a classical test case for compressible Euler equations, typically used to assess the PEP property. In our framework, however, the density-wave test has been chosen as a simple one-dimensional case in which the entropy-conservation properties of the novel formulation can be assessed with minimum occurrences of locally uniform temperature distributions. This allows us to test the original fluxes by minimizing the use of the alternative non-singular formulations. In fact, for this particular test-case, the threshold value for switching to a non-singular formulation is set to $\delta^{\,+}T<10^{-14}$.
@@
with $n=1$ and $u_{0}=10\,\text{m/s}$. We define a characteristic time $t_{\text{ref}}=L/u_{0}=0.1\,\text{s}$, indicating the time taken by the initial condition to travel across the whole domain.
@@
Fig. 2 shows the results of simulations with final time $t=0.6$ s and $N=40$ intervals, with a Courant number set to $0.005$ for the Van der Waals (2nd order, Fig. 2(a)) case and $0.002$ for the Peng-Robinson (4th order, Fig. 2(b)) case, in order to be able to neglect time-integration errors. The performances of the novel fluxes are compared with those of the Ranocha’s and AEC(0) formulations.
The insets clearly show that the new formulation reduces the global error on entropy production up to the machine precision, whereas the classical formulations exhibit a non-negligible spurious production of entropy which accumulates in time. As it is possible to note from Fig. 2, the Ranocha scheme seems to perform better than the AEC(0) scheme. The same simulations, not included here, have also been carried out with the AEC scheme with two additional terms of the series expansion which has been chosen as the non-singular flux to be used locally in the presence of singularities for the new EC flux. The results showed an essentially identical behavior to that of the scheme of Ranocha regarding the error $\langle\rho s\rangle$, which is in agreement with the theory presented in [30].
@@ 7.2 Transcritical double shear layer
A two-dimensional double shear layer is simulated in the rectangle $L_{x},L_{x}]\times[-L_{y},L_{y}]$, with $L_{x}=0.5\,\text{m}$ and $L_{y}=0.25\,\text{m}$, discretized in $N_{x}\times N_{y}=32\times 16$ intervals on a uniform mesh with periodic boundary conditions and initial conditions
@@
where $A=B=3/8$, $\epsilon=0.1$, $\delta=1/15$ and $k=3$. Courant number is set to $0.005$. The constants $u_{0}=20\,\text{m/s}$ and $T_{0}=110\,\text{K}$ are the streamwise component of the velocity and the temperature on the center line $(x,0,0)$, respectively. The $y$- component of velocity, $v$, is initialized to provide a perturbation that triggers three roll-up vortices, while exponentially decaying with the distance from the centerline. Following [18], we set a bulk pressure $p=2p_{c}$, whereas the density $\rho$ is evaluated from the equation of state.
@@
From Fig. 3, reporting the entropy production for the various schemes, it is seen that the theoretical conservation properties of the new formulation are confirmed also in this test. In this case, the harmonic numerical flux on internal energy provides a smaller error on the conservation of $\rho s$ if compared to the Ranocha scheme. In particular, the difference between the two discretizations appears to be more relevant for the Peng-Robinson equation of state. The threshold value for this test-case has been set to $\delta^{\,+}T<10^{-3}$.
7.3 Supercritical Taylor-Green vortices
The 3D Taylor-Green Vortex (TGV) test case is simulated in the tri-periodic box of size $L_{x}\times L_{y}\times L_{z}=(2\pi L)^{3}$, with $L=1\,\text{m}$, discretized in $N=32^{3}$ evenly-spaced points. The threshold value on temperature is set to be $\delta^{\,+}T<10^{-3}$. The simulations have been carried out in the open-source, scalable, GPU-accelerated code STREAmS-2.0 [57, 58], which performs high-fidelity direct numerical simulations of high-speed compressible flows. The original code, which discretizes the compressible Navier-Stokes equations with the perfect gas EoS, has been adapted by including the Van Der Waals and Peng-Robinson models, as well as the newly derived fluxes for the convective terms given in Eq. (50). The spatial integration is performed with a sixth-order central scheme, while the third-order, low-storage Wray’s RK method already included in the original version of STREAmS-2.0, is used for the time-integration, with a $\text{CFL}=0.001$ based on the speed of sound.
The initial conditions are set as
$\begin{cases}\rho(x,y,z,0)&=\rho_{0}\\ u(x,y,z,0)&=u_{0}\sin{x}\cos{y}\cos{z}\\ v(x,y,z,0)&=-u_{0}\cos{x}\sin{y}\cos{z}\\ w(x,y,z,0)&=0\\ p(x,y,z,0)&=p_{0}+\dfrac{\rho_{0}u_{0}^{2}}{16}(\cos{2x}+\cos{2y})(2+\cos{2z})\end{cases}$ (62)
with $u_{0}=20\,\text{m/s}$, $\rho_{0}=0.8\rho_{c}$, $p_{0}=10\,\text{MPa}$. The reference time $t_{\text{ref}}$ is defined as $L/u_{0}=0.05\,\text{s}$. The resulting temperature is $T\approx 185\,\text{K}$, so that the fluid is in supercritical conditions, and the subsequent Mach number is $M\approx 0.05$.
Figure 4: Entropy conservation and thermodynamic fluctuations for the 3D TGV with the Van Der Waals equation of state. Colors are set as in the previous figures, while green lines with circles represent the KEEP scheme, which is the built-in spatial discretization in STREAmS-2.0.
In Fig. 4 and 5 the entropy-conservation and the temperature fluctuations are shown for both the Van Der Waals and Peng-Robinson equations of state, respectively. In addition to the Ranocha and AEC(0) scheme already used in the previous test cases, we simulated the present test case by also using the KEEP schemes [31], already implemented in the STREAmS-2.0 code. The proposed formulation consistently shows near machine-zero entropy conservation properties, although a slightly drifting behavior, which is attributed to the low-storage implementation of the third-order RK method, is visible for both the numerical tests. All the compared discretizations show that the fluctuations and the error on entropy conservation stabilize around an almost steady state after a transient behavior.
Figure 5: Entropy conservation and thermodynamic fluctuations for the 3D TGV with the Peng-Robinson equation of state. Colors are set as in the previous figures, while green lines with circles represent the KEEP scheme, which is the built-in spatial discretization in STREAmS-2.0.
Density fluctuations (not shown here) have also been obtained, giving the same information provided by the temperature fluctuations.
@@ 8 Conclusions
The derived formulation is, to the best of our knowledge, the first entropy-conserving discretization of the compressible Euler equations specifically designed for real gases. Its structure-preserving properties make it an ideal candidate for high-fidelity simulations of turbulent flows where an arbitrary equation of state is used. Moreover, it constitutes the first building block toward a fully structure-preserving discretization of compressible flow equations for real gas applications. In fact, the condition introduced in this work on the discretization of the convective term of internal energy introduces a whole family of EC schemes for which, however, no constraint is imposed on the choice of the numerical mass flux. The use of this additional degree of freedom could be exploited to impart the scheme with supplementary structure-preserving properties, an example of which being the pressure equilibrium preservation property that is often desirable in the contexts of multi-component gases or multi-phase flows. The exploration of this topic is left for future research.
Acknowledgments
We acknowledge the CINECA award under the ISCRA initiative, for the availability of high-performance computing resources and support.